# RovoDev Code Reviewer: A Large-Scale Online Evaluation of LLM-based Code Review Automation at Atlassian

### v1

@@ Abstract.
Large Language Models (LLMs)-powered code review automation has the potential to transform code review workflows. Despite the advances of LLM-powered code review comment generation approaches, several practical challenges remain for designing enterprise-grade code review automation tools. In particular, this paper aims at answering the practical question: how can we design a review-guided, context-aware, quality-checked code review comment generation without fine-tuning? In this paper, we present RovoDev Code Reviewer, an enterprise-grade LLM-based code review automation tool designed and deployed at scale within Atlassian’s development ecosystem with seamless integration into Atlassian’s Bitbucket. Through the offline, online, user feedback evaluations over a one-year period, we conclude that RovoDev Code Reviewer is (1) effective in generating code review comments that could lead to code resolution for 38.70% (i.e., comments that triggered code changes in the subsequent commits); and (2) offers the promise of accelerating feedback cycles (i.e., decreasing the PR cycle time by 30.8%), alleviating reviewer workload (i.e., reducing the number of human-written comments by 35.6%), and improving overall software quality (i.e., finding errors with actionable suggestions).
@@
conference: the 48th International Conference on Software Engineering (ICSE); April 12–18, 2026; Rio De Janeiro, Brazil.ccs: Software and its engineering Software development techniques
@@ 1. Introduction
Code review is a cornerstone of modern software engineering [bacchelli2013expectations, rigby2011broadcast, rigby2013decade, macleod2017code], serving as a critical quality assurance practice that helps teams identify defects, share knowledge, and maintain high coding standards. However, as software projects grow in size and complexity, manual code review becomes increasingly time-consuming and resource-intensive [thongtanunam2015should], often leading to bottlenecks in the development process. Automating aspects of code review using advances in Large Language Models (LLMs) [tufano2025automating, tufano2024code, tufano2019learning, thongtanunam2014improving, thongtanunam2022autotransform, hong2022commentfinder, hong2022should] could speed up code review processes, particularly, code review comment generation, defined as a generative task to generate code review comments written in natural languages for a given code change [hong2022commentfinder, frommgen2024resolving, hong2025retrieval, olewicki2024impact, tufano2022using].
Despite these advances, several practical challenges remain for designing enterprise-grade code review automation tools. First, data privacy and security are paramount, especially when processing sensitive customer code and metadata, making it infeasible to fine-tune LLMs on proprietary or user-generated content in many industrial contexts. Second, review guidelines play a critical role to guide inexperienced reviewers to conduct a code review, yet it remains largely ignored in the recent LLM-powered code review comment generation approaches [olewicki2024impact]. Third, most retrieval-augmented generation (RAG) approaches [olewicki2024impact, hong2025retrieval] rely on rich historical data, which may not be available for newly created or context-limited projects. Finally, LLMs are prone to generating noisy or hallucinated comments that may be vague, non-actionable, or factually incorrect [liu2025too, she2023pitfalls], potentially diminishing the utility of automated code review tools.
@@ 2.1. Modern Code Review at Atlassian
Modern code review (MCR) has become a cornerstone of contemporary software engineering practice [rigby2011broadcast, rigby2013decade], replacing traditional, formal inspections with lightweight, tool-supported, and asynchronous processes. Unlike earlier approaches that relied on in-person meetings and checklists, MCR is typically conducted using platforms such as Bitbucket [bitbucket], Gerrit [gerrit], GitHub [github], GitLab [gitlab], which facilitate distributed and collaborative review of code changes [bacchelli2013expectations].
At Atlassian (a global technology company known for its suite of software development tools, such as, Bitbucket, Jira, and Confluence) [atlassian], code review is a mandatory and deeply embedded practice across all software engineering teams. Every code change, regardless of size, is subject to peer review before it can be merged into the main codebase, ensuring its code quality, maintainability, and security. This process is facilitated primarily through pull requests in Bitbucket , where software engineers submit their changes and request feedback from reviewers. The process is collaborative, with reviewers and authors engaging in discussions to clarify intent, resolve issues, and reach consensus before changes are approved.
@@ 2.2. Large Language Models-based Code Review Automation and Limitations
Code review automation [tufano2025automating] has been an active area of research, driven by the need to improve software quality, engineers’ productivity, and the code review speed. With the rise of machine learning and natural language processing, researchers have explored data-driven approaches to code review automation [li2022automating]. For example, reviewer recommendation [balachandran2013reducing, thongtanunam2015should, thongtanunam2014improving, hannebauer2016automatically], code review prioritization [fu2022linevul, pornprasit2021jitline, wattanakriengkrai2020predicting], line-level localization [hong2022should], code review comment generation [hong2025retrieval, frommgen2024resolving, olewicki2024impact, liu2025too] and recommendation [hong2022commentfinder], code review usefulness [bosu2015characteristics, pangsakulyanont2014assessing], and code refinement [thongtanunam2022autotransform, tufan2021towards, tufano2019learning, fu2022VulRepair, lu2023llama].
Figure 2. An overview of our RovoDev Code Reviewer.
@@
LLM-driven code review comment generation is an emerging research area that leverages Large Language Models (LLMs) to automatically generate review comments for a given code change. Recent studies have explored the use of off-the-shelf LLMs, such as, Neural Machine Translation [tufan2021towards], T5 [tufano2022using], CodeT5 [tufano2024code], Llama [llamareviewer], GPT-3.5 [pornprasit2024fine], and GPT-4o [olewicki2024impact]. Many studies have demonstrated that fine-tuning LLMs on task-specific data yields the best effectiveness for code review comment generation. This process typically relies on datasets consisting of code and corresponding human-written comments, often sourced from publicly available repositories (e.g., TufanoT5 [tufan2021towards]). However, several practical yet unsolved challenges arise when designing an enterprise-grade code review comment generation tool.
@@
First, an enterprise-grade code review comment generation tool should protect customers’ code and related metadata. For Atlassian’s enterprise customers, data privacy and security are top concerns—nearly half cite these as critical blockers to adopting AI-powered software development tools [torgerson2024datasecurity]. Mishandling customer data can lead to severe consequences [atlassian2025privacy], including breaches of confidentiality, loss of customer trust, regulatory non-compliance (such as GDPR or CCPA), and significant financial or reputation damage. As a result, fine-tuning large language models, which is common in the existing work [hong2025retrieval, tufano2021towards], for code review comment generation may pose a risk of data security and privacy for enterprise-grade automated code review tools.
@@
Second, an enterprise-grade code review comment generation tool should be equipped with high-standard code review guidelines. Code review guidelines or checklists often serve as a fundamental component of code review processes. Prior study found that checklist-based code review is an effective mechanism for helping inexperienced reviewers to conduct code review processes, since the checklist can guide them to focus on the important issues or defects when performing code reviews [chong2021assessing]. However, such code review checklists have been ignored in the recent automated code review approaches [olewicki2024impact].
@@
Third, an enterprise-grade code review comment generation tool should be context-aware, while being capable of supporting newly created software projects with limited contextual information. Recent work proposed to use Retrieval-Augmented Generation (RAG) [hong2025retrieval, olewicki2024impact] to retrieve within-repository contextual information to further improve the effectiveness of code review comment generation task. For example, Hong et al. [hong2025retrieval] proposed to retrieve review comments corresponding to code snippets that are most similar to the given input code. Similarly, Olewick et al. [olewicki2024impact] proposed to retrieve related comment examples for a given code change. These RAG-based approaches are typically designed for mature, long-standing software projects with substantial historical data. However, they are less suitable for newly created projects that lack sufficient historical context [pornprasit2024fine].
@@
Finally, an enterprise-grade code review comment generation should generate valid and factually correct code review comments. Defined by Tufano et al. [tufano2024code], valid review comments refer to comments that provide clear suggestions aimed at improving the source code quality. Liu et al. [liu2025too] found that many code review comments could be noisy comments (e.g., vague or non-actionable feedback). For example, “Add a blank line here”. Such noisy comments may lead LLMs to generate low-quality review comments. To make matters worse, LLMs are prone to hallucination [ji2023survey, zhang2025llm, huang2025survey]—i.e., generating code review comments that are inaccurate, inconsistent, or nonsensical based on the provided code change. Zhang et al. [zhang2025llm] found that such hallucination is commonly found in the code generation task. Such invalid code review comments may diminish the overall utility of the enterprise-grade automated code review tools. These practical challenges of designing an enterprise-graded code review comment generation lead us to formulate the ultimate research question: How can we design a review-guided, context-aware, quality-checked code review comment generation without fine-tuning?
@@ 3. Atlassian’s RovoDev Code Reviewer
A Comment Quality Check Component on Factual Correctness. To ensure that code review comments being generated are not hallucinated (i.e., aligns with a given code change and its associated pull request), we use a LLM-as-a-Judge [zheng2023judging] approach to verify if a given comment is aligned with a given pull request and the area of code change or not.
@@ 3.1. A Context-Aware Review-Guided Comment Generation
Contextual information plays an important role in many tasks in software engineering [xu2020towards, wen2018context], especially, code review automation [pornprasit2025context]. Prior studies leverage various types of contextual information e.g., similar code changes to augment the code review comment generation task [hong2025retrieval, olewicki2024impact, hong2022commentfinder]. However, such similar code changes require information access to within- or across-repositories, which may pose risks of data privacy and security and may not be available for the newly-created or context-limited software projects.
@@
$\bullet$ Persona ($P$) is an explicit instruction to set the LLM to act as a software engineer, similar to prior studies [olewicki2024impact, pornprasit2024fine].
@@ 3.2. Comment Quality Check on Factual Correctness
It is possible that the set of generated code review comments ($R$) may be inaccurate, inconsistent, or nonsensical based on the provided code change, due to the LLM hallucination [ji2023survey, zhang2025llm, huang2025survey]. To avoid LLM being hallucinated, we use LLM-as-a-Judge [li2024generation, zheng2023judging] ($J_{\mathrm{fact}}$) to evaluate if the set of generated code review comments ($R$) are factually correct (i.e., the generated comment is aligned with a given code change). Our LLM Judge ($J_{\mathrm{fact}}$) is based on a prompting technique on a gpt-4o-mini model where our prompt consists of persona, task definition, instructions, and assessment guidelines. Formally speaking, given an LLM Judge $J_{\mathrm{fact}}$, the assessment process can be formulated as $R=\{C_{i},...,C_{j}\}$, where $C_{i}$ is the $i^{th}$ candidate to be judged and $R$ is the judging result. We utilize a selection-based judgment methodology [li2024generation] that returns a binary outcome (True/False) in order to select one or more valid comment candidates to avoid the subjective threshold selection that is commonly presented in the score-based and ranking-based judgment. Therefore, only code review comments that are factually correct will be proceed to the next component.
@@ 3.3. Comment Quality Check on Actionability
Although the generated-code review comments are factually correct, they may be vague (e.g., “Needs improvement”), nitpicking without context (e.g., “Add a blank line here”), unfocused or off-topic comments (e.g., “I saw something similar in another project”), comments lacking specificity (e.g., “Is this the best way?”), and non-actionable comments (e.g., “Good job!”). To address this challenge, we develop a ModernBERT-based Comment Quality Gate model to filter out such noisy code review comments generated by LLMs. ModernBERT [warner2024smarter] is a modernized version of the encoder-only BERT model trained on 2 trillion tokens with a native 8,192 sequence length, allowing a longer context length for our Comment Quality Check component to infer if a given comment is of high quality or not. To this end, we develop our Comment Quality Check model to capture the characteristics of RovoDev-generated code review comments that are most likely to drive code resolution. We fine-tuned our ModernBERT model using five months of over 50,000 high-quality RovoDev-generated code review comments from Atlassian’s proprietary repositories. Note that we only use the historical data for model training to avoid data leakage in the online experiment. Each data point consists of a RovoDev-generated comment paired with a label indicating whether it led to a code resolution (<<comment, resolved?>>). Therefore, only code review comments that are valid and most likely to drive code resolution will be recommended to the given PR.
@@ 3.4. A Seamless Integration of RovoDev Code Reviewer into Atlassian Bitbucket
Following our prior study [takerngsaksiri2024human] and Atlassian’s Human-AI Collaboration Strategy as the Future of Teamwork, a human-in-the-loop paradigm is maintained. Rather than automating the whole code review process, RovoDev Code Reviewer serves as a code review assistant, aiming to generate review comments and suggestions, but final decisions remain with human reviewers, supporting trust, accountability, and iterative improvement, and respecting human autonomy. To ensure the scalability of our RovoDev Code Reviewer, we use an Event-driven Architecture (EDA), i.e., a software design paradigm that determines the flow of RovoDev Code Reviewer system by events changes. Briefly speaking, when a pull request is created, RovoDev is triggered to begin the automated code review process as follows:
@@ 4.2. A Large-Scale Online Evaluation
Motivation. As is common in code review automation research, any offline evaluation often relies on human-written comments as ground truth [hong2025retrieval, frommgen2024resolving, olewicki2024impact, liu2025too, tufan2021towards, tufano2022using, tufano2024code, llamareviewer, pornprasit2024fine] with several inherent limitations. First, many code review benchmark dataset is static [tufan2021towards, tufano2022using, tufano2024code, llamareviewer], which does not capture the full complexity, variability, or evolving nature of real-world software engineering tasks. Second, the commonly-used semantic matching evaluation metrics (e.g., an exact match, BLEU, semantic similarity) [liu2025too] is constrained by the similarity to the human-written comments, but does not adequately reflect the practical usefulness (i.e., whether the comment actually facilitates code resolution). In reality, code review comments can be written differently, and can be uniquely different from the human-written comments, yet share the same or similar intents. Therefore, an online evaluation is critically needed to rigorously investigate the practical value of code review comments generated by RovoDev Code Reviewer.
@@
Results.  RovoDev Code Reviewer achieves a code resolution rate of 38.7%, which is 12.9% less than the code resolution rate of human-written comments (44.45%). Figure 3(a) presents the Code Resolution Rate (CRR) of RovoDev-generated comments and human-written comments over a one-year period (Aug 2024 to July 2025), using a rolling average of 7-day (i.e., an average CRR value within a moving window of 7-days). This one-year period includes the beginning of RovoDev’s internal dogfooding (Aug 2025) to the widespread enablement, until July 15 2025. When considering the recent one-week period between July 10 2025 and July 15 2025, RovoDev Code Reviewer achieves a code resolution rate of 38.70%, while human-written comments achieves a code resolution rate of 44.45%. This leads to a maximum difference of 5.8 percentage points, or a relative percent difference of just 12.9% ($\frac{38.70-44.45}{44.45}$) from the code resolution rate of human-written comments, highlighting the high practical value and quality of RovoDev-generated code review comments. In addition, the most resolved types of code review comments are related to readability, bugs, and maintainability [goldman2025types].
@@
Motivation. User evaluation feedback is a critical component in both academic research and product development, as it provides direct insights into user needs, behaviors, and experiences. However, there exist qualitative research methods such as survey and interview used for evaluating automated code review in practice. For example, Cihan et al. [cihan2024automated] conducted a board survey with 22 practitioners to investigate the practitioner’s perception on automated code review in practice. However, such survey studies only capture the general opinions from software engineers rather than the specific fine-grained qualitative feedback at each individual comment (i.e., why a particular code review comment is useful or not useful).
Research Methodology. To address the gap, we collected feedback at each individual comment. For each comment generated by RovoDev, we explicitly ask software engineers to rate the comment quality using the thumbs-up or thumbs-down. However, similar to a case study at Google [vijayvergiya2024ai], we observe that software engineers rarely clicked the thumbs up/thumbs down buttons. Therefore, we also ask the software engineers to provide an explicit feedback if they find a RovoDev-generated comment particularly useful or not with an option for our RovoDev team to make a direct contact with the software engineers or not, respecting user privacy. All of the comments were collected and fed into the Atlassian’s Jira Service Management Cloud System for feedback analysis, prioritization, and management. We conducted a thematic analysis to identify the common themes of the feedback data using an open-coding approach. Following a Reflexive Thematic Analysis (RTA) [byrne2022worked], we manually reviewed each feedback and assigned a code and theme that is related to the feedback. With the reflexive approach, codes and themes will be evolved throughout the process, meaning that there is no fixed codebook and boundaries of codes can be redrawn, split, merged, or promoted to themes as understanding deepens. For example, “This is an amazing catch” and “Nope we were wrong, the AI was right!” were grouped into one theme.

### v2

@@ Abstract.
Large Language Models (LLMs)-powered code review automation has the potential to transform code review workflows. Despite the advances of LLM-powered code review comment generation approaches, several practical challenges remain for designing enterprise-grade code review automation tools. In particular, this paper aims at answering the practical question: how can we design a review-guided, context-aware, quality-checked code review comment generation without fine-tuning? In this paper, we present RovoDev Code Reviewer, an enterprise-grade LLM-based code review automation tool designed and deployed at scale within Atlassian’s development ecosystem with seamless integration into Atlassian’s Bitbucket. Through the offline, online, user feedback evaluations over a one-year period, we conclude that RovoDev Code Reviewer is effective in generating code review comments that could lead to code resolution for 38.70% (i.e., comments that triggered code changes in the subsequent commits); and offers the promise of accelerating feedback cycles (i.e., decreasing the PR cycle time by 30.8%), alleviating reviewer workload (i.e., reducing the number of human-written comments by 35.6%), and improving overall software quality (i.e., finding errors with actionable suggestions).
@@
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIP ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786583.3786851isbn: 979-8-4007-2426-8/2026/04ccs: Software and its engineering Software development techniques
@@ 1. Introduction
Code review is a cornerstone of modern software engineering [2, 36, 37, 29], serving as a critical quality assurance practice that helps teams identify defects, share knowledge, and maintain high coding standards. However, as software projects grow in size and complexity, manual code review becomes increasingly time-consuming and resource-intensive [43], often leading to bottlenecks in the development process. Automating aspects of code review using advances in Large Language Models (LLMs) [46, 47, 45, 41, 42, 20, 21] could speed up code review processes, particularly, code review comment generation, defined as a generative task to generate code review comments written in natural languages for a given code change [20, 11, 19, 31, 48].
Despite these advances, several practical challenges remain for designing enterprise-grade code review automation tools. First, data privacy and security are paramount, especially when processing sensitive customer code and metadata, making it infeasible to fine-tune LLMs on proprietary or user-generated content in many industrial contexts. Second, review guidelines play a critical role to guide inexperienced reviewers to conduct a code review, yet it remains largely ignored in the recent LLM-powered code review comment generation approaches [31]. Third, most retrieval-augmented generation (RAG) approaches [31, 19] rely on rich historical data, which may not be available for newly created or context-limited projects. Finally, LLMs are prone to generating noisy or hallucinated comments that may be vague, non-actionable, or factually incorrect [26, 38], potentially diminishing the utility of automated code review tools.
@@ 2.1. Modern Code Review at Atlassian
Modern code review (MCR) has become a cornerstone of contemporary software engineering practice [36, 37], replacing traditional, formal inspections with lightweight, tool-supported, and asynchronous processes. Unlike earlier approaches that relied on in-person meetings and checklists, MCR is typically conducted using platforms such as Bitbucket [4], Gerrit [14], GitHub [15], GitLab [40], which facilitate distributed and collaborative review of code changes [2].
At Atlassian (a global technology company known for its suite of software development tools, such as, Bitbucket, Jira, and Confluence) [10], code review is a mandatory and deeply embedded practice across all software engineering teams. Every code change, regardless of size, is subject to peer review before it can be merged into the main codebase, ensuring its code quality, maintainability, and security. This process is facilitated primarily through pull requests in Bitbucket , where software engineers submit their changes and request feedback from reviewers. The process is collaborative, with reviewers and authors engaging in discussions to clarify intent, resolve issues, and reach consensus before changes are approved.
@@ 2.2. Large Language Models-based Code Review Automation and Limitations
Code review automation [46] has been an active area of research, driven by the need to improve software quality, engineers’ productivity, and the code review speed. With the rise of machine learning and natural language processing, researchers have explored data-driven approaches to code review automation [25]. For example, reviewer recommendation [3, 43, 41, 18], code review prioritization [13, 33, 53], line-level localization [21], code review comment generation [19, 11, 31, 26] and recommendation [20], code review usefulness [5, 32], and code refinement [42, 50, 45, 12, 28].
@@
LLM-driven code review comment generation is an emerging research area that leverages Large Language Models (LLMs) to automatically generate review comments for a given code change. Recent studies have explored the use of off-the-shelf LLMs, such as, Neural Machine Translation [50], T5 [48], CodeT5 [47], Llama [27], GPT-3.5 [34], and GPT-4o [31]. Many studies have demonstrated that fine-tuning LLMs on task-specific data yields the best effectiveness for code review comment generation. This process typically relies on datasets consisting of code and corresponding human-written comments, often sourced from publicly available repositories (e.g., TufanoT5 [50]). However, several practical yet unsolved challenges arise when designing an enterprise-grade code review comment generation tool.
@@
First, an enterprise-grade code review comment generation tool should protect customers’ code and related metadata. For Atlassian’s enterprise customers, data privacy and security are top concerns—nearly half cite these as critical blockers to adopting AI-powered software development tools [44]. Mishandling customer data can lead to severe consequences [1], including breaches of confidentiality, loss of customer trust, regulatory non-compliance (such as GDPR or CCPA), and significant financial or reputation damage. As a result, fine-tuning large language models, which is common in the existing work [19, 49], for code review comment generation may pose a risk of data security and privacy for enterprise-grade automated code review tools.
@@
Second, an enterprise-grade code review comment generation tool should be equipped with high-standard code review guidelines. Code review guidelines or checklists often serve as a fundamental component of code review processes. Prior study found that checklist-based code review is an effective mechanism for helping inexperienced reviewers to conduct code review processes, since the checklist can guide them to focus on the important issues or defects when performing code reviews [8]. However, such code review checklists have been ignored in the recent automated code review approaches [31].
@@
Third, an enterprise-grade code review comment generation tool should be context-aware, while being capable of supporting newly created software projects with limited contextual information. Recent work proposed to use Retrieval-Augmented Generation (RAG) [19, 31] to retrieve within-repository contextual information to further improve the effectiveness of code review comment generation task. For example, Hong et al. [19] proposed to retrieve review comments corresponding to code snippets that are most similar to the given input code. Similarly, Olewick et al. [31] proposed to retrieve related comment examples for a given code change. These RAG-based approaches are typically designed for mature, long-standing software projects with substantial historical data. However, they are less suitable for newly created projects that lack sufficient historical context [34].
@@
Finally, an enterprise-grade code review comment generation should generate valid and factually correct code review comments. Defined by Tufano et al. [47], valid review comments refer to comments that provide clear suggestions aimed at improving the source code quality. Liu et al. [26] found that many code review comments could be noisy comments (e.g., vague or non-actionable feedback). For example, “Add a blank line here”. Such noisy comments may lead LLMs to generate low-quality review comments. To make matters worse, LLMs are prone to hallucination [23, 56, 22]—i.e., generating code review comments that are inaccurate, inconsistent, or nonsensical based on the provided code change. Zhang et al. [56] found that such hallucination is commonly found in the code generation task. Such invalid code review comments may diminish the overall utility of the enterprise-grade automated code review tools. These practical challenges of designing an enterprise-graded code review comment generation lead us to formulate the ultimate research question: How can we design a review-guided, context-aware, quality-checked code review comment generation without fine-tuning?
@@ 3. Atlassian’s RovoDev Code Reviewer
A Comment Quality Check Component on Factual Correctness. To ensure that code review comments being generated are not hallucinated (i.e., aligns with a given code change and its associated pull request), we use a LLM-as-a-Judge [57] approach to verify if a given comment is aligned with a given pull request and the area of code change or not.
@@ 3.1. A Context-Aware Review-Guided Comment Generation
Contextual information plays an important role in many tasks in software engineering [55, 54], especially, code review automation [35]. Prior studies leverage various types of contextual information e.g., similar code changes to augment the code review comment generation task [19, 31, 20]. However, such similar code changes require information access to within- or across-repositories, which may pose risks of data privacy and security and may not be available for the newly-created or context-limited software projects.
@@
$\bullet$ Persona ($P$) is an explicit instruction to set the LLM to act as a software engineer, similar to prior studies [31, 34].
@@ 3.2. Comment Quality Check on Factual Correctness
It is possible that the set of generated code review comments ($R$) may be inaccurate, inconsistent, or nonsensical based on the provided code change, due to the LLM hallucination [23, 56, 22]. To avoid LLM being hallucinated, we use LLM-as-a-Judge [24, 57] ($J_{\mathrm{fact}}$) to evaluate if the set of generated code review comments ($R$) are factually correct (i.e., the generated comment is aligned with a given code change). Our LLM Judge ($J_{\mathrm{fact}}$) is based on a prompting technique on a gpt-4o-mini model where our prompt consists of persona, task definition, instructions, and assessment guidelines. Formally speaking, given an LLM Judge $J_{\mathrm{fact}}$, the assessment process can be formulated as $R=\{C_{i},...,C_{j}\}$, where $C_{i}$ is the $i^{th}$ candidate to be judged and $R$ is the judging result. We utilize a selection-based judgment methodology [24] that returns a binary outcome (True/False) in order to select one or more valid comment candidates to avoid the subjective threshold selection that is commonly presented in the score-based and ranking-based judgment. Therefore, only code review comments that are factually correct will be proceed to the next component.
@@ 3.3. Comment Quality Check on Actionability
Although the generated-code review comments are factually correct, they may be vague (e.g., “Needs improvement”), nitpicking without context (e.g., “Add a blank line here”), unfocused or off-topic comments (e.g., “I saw something similar in another project”), comments lacking specificity (e.g., “Is this the best way?”), and non-actionable comments (e.g., “Good job!”). To address this challenge, we develop a ModernBERT-based Comment Quality Gate model to filter out such noisy code review comments generated by LLMs. ModernBERT [52] is a modernized version of the encoder-only BERT model trained on 2 trillion tokens with a native 8,192 sequence length, allowing a longer context length for our Comment Quality Check component to infer if a given comment is of high quality or not. To this end, we develop our Comment Quality Check model to capture the characteristics of RovoDev-generated code review comments that are most likely to drive code resolution. We fine-tuned our ModernBERT model using five months of over 50,000 high-quality RovoDev-generated code review comments from Atlassian’s proprietary repositories. Note that we only use the historical data for model training to avoid data leakage in the online experiment. Each data point consists of a RovoDev-generated comment paired with a label indicating whether it led to a code resolution (<<comment, resolved?>>). Therefore, only code review comments that are valid and most likely to drive code resolution will be recommended to the given PR.
@@ 3.4. A Seamless Integration of RovoDev Code Reviewer into Atlassian Bitbucket
Following our prior study [39] and Atlassian’s Human-AI Collaboration Strategy as the Future of Teamwork, a human-in-the-loop paradigm is maintained. Rather than automating the whole code review process, RovoDev Code Reviewer serves as a code review assistant, aiming to generate review comments and suggestions, but final decisions remain with human reviewers, supporting trust, accountability, and iterative improvement, and respecting human autonomy. To ensure the scalability of our RovoDev Code Reviewer, we use an Event-driven Architecture (EDA), i.e., a software design paradigm that determines the flow of RovoDev Code Reviewer system by events changes. Briefly speaking, when a pull request is created, RovoDev is triggered to begin the automated code review process as follows:
@@ 4.2. A Large-Scale Online Evaluation
Motivation. As is common in code review automation research, any offline evaluation often relies on human-written comments as ground truth [19, 11, 31, 26, 50, 48, 47, 27, 34] with several inherent limitations. First, many code review benchmark dataset is static [50, 48, 47, 27], which does not capture the full complexity, variability, or evolving nature of real-world software engineering tasks. Second, the commonly-used semantic matching evaluation metrics (e.g., an exact match, BLEU, semantic similarity) [26] is constrained by the similarity to the human-written comments, but does not adequately reflect the practical usefulness (i.e., whether the comment actually facilitates code resolution). In reality, code review comments can be written differently, and can be uniquely different from the human-written comments, yet share the same or similar intents. Therefore, an online evaluation is critically needed to rigorously investigate the practical value of code review comments generated by RovoDev Code Reviewer.
@@
Results.  RovoDev Code Reviewer achieves a code resolution rate of 38.7%, which is 12.9% less than the code resolution rate of human-written comments (44.45%). Figure 3(a) presents the Code Resolution Rate (CRR) of RovoDev-generated comments and human-written comments over a one-year period (Aug 2024 to July 2025), using a rolling average of 7-day (i.e., an average CRR value within a moving window of 7-days). This one-year period includes the beginning of RovoDev’s internal dogfooding (Aug 2025) to the widespread enablement, until July 15 2025. When considering the recent one-week period between July 10 2025 and July 15 2025, RovoDev Code Reviewer achieves a code resolution rate of 38.70%, while human-written comments achieves a code resolution rate of 44.45%. This leads to a maximum difference of 5.8 percentage points, or a relative percent difference of just 12.9% ($\frac{38.70-44.45}{44.45}$) from the code resolution rate of human-written comments, highlighting the high practical value and quality of RovoDev-generated code review comments. In addition, the most resolved types of code review comments are related to readability, bugs, and maintainability [16].
@@
Motivation. User evaluation feedback is a critical component in both academic research and product development, as it provides direct insights into user needs, behaviors, and experiences. However, there exist qualitative research methods such as survey and interview used for evaluating automated code review in practice. For example, Cihan et al. [9] conducted a board survey with 22 practitioners to investigate the practitioner’s perception on automated code review in practice. However, such survey studies only capture the general opinions from software engineers rather than the specific fine-grained qualitative feedback at each individual comment (i.e., why a particular code review comment is useful or not useful).
Research Methodology. To address the gap, we collected feedback at each individual comment. For each comment generated by RovoDev, we explicitly ask software engineers to rate the comment quality using the thumbs-up or thumbs-down. However, similar to a case study at Google [51], we observe that software engineers rarely clicked the thumbs up/thumbs down buttons. Therefore, we also ask the software engineers to provide an explicit feedback if they find a RovoDev-generated comment particularly useful or not with an option for our RovoDev team to make a direct contact with the software engineers or not, respecting user privacy. All of the comments were collected and fed into the Atlassian’s Jira Service Management Cloud System for feedback analysis, prioritization, and management. We conducted a thematic analysis to identify the common themes of the feedback data using an open-coding approach. Following a Reflexive Thematic Analysis (RTA) [6], we manually reviewed each feedback and assigned a code and theme that is related to the feedback. With the reflexive approach, codes and themes will be evolved throughout the process, meaning that there is no fixed codebook and boundaries of codes can be redrawn, split, merged, or promoted to themes as understanding deepens. For example, “This is an amazing catch” and “Nope we were wrong, the AI was right!” were grouped into one theme.